# FC-KAN: Function Combinations in Kolmogorov-Arnold Networks

### v1

@@ 4.2 Model Performance
4.3 Comparison of Combination Types
While employing a quadratic function representation for the output of FC-KAN, we are also interested in exploring how different output combination methods affect model performance. In this experiment, we use FC-KAN (DoG+BS) with several output combination methods: sum, element-wise product, addition of sum and element-wise product, quadratic function representation, and concatenation.
@@ 4.3 Comparison of Combination Types
Dataset Combined Type Train. Acc. Val. Acc. F1 Time (seconds)
@@
From the results in Table 3, the quadratic function best represents the combination output and outperforms other combinations, although its models require more training time. Meanwhile, the output combination by concatenation shows the worst results. In MNIST, the addition of sum and element-wise product demonstrates very competitive performance while requiring the least training time. Except for concatenation, the other four combinations can easily achieve 100% training accuracy. In Fashion-MNIST, the element-wise product combination is only surpassed by the quadratic function, but it takes 6.3% less training time.
@@ 6 Conclusion
We introduced FC-KAN, which uses various popular mathematical functions to represent data features and combines their outputs using different methods, primarily through element-wise operations in low-dimensional layers, to address image classification problems. In the experiments, we designed FC-KAN to combine pairs of functions, such as B-splines, wavelets, and radial basis functions, using several output combinations on the MNIST and Fashion-MNIST datasets. We found that FC-KAN outperformed MLP and other KAN models with the same network structure, although it required more training time. Among the variants, FC-KAN (DoG+BS), which combines the Difference of Gaussians and B-splines with a quadratic function representation in the output, achieved the best results on both datasets. FC-KAN demonstrates promising potential in using function combinations to design KANs and improve model performance. However, there are several aspects, such as exploring other functions and combinations to extract data features for further model enhancement, that we would like to investigate. These will be the focus of our future work.

### v2

@@ 4.2 Model Performance
4.3 Comparison of Combination Methods
While employing a quadratic function representation for the output of FC-KAN, we are also interested in exploring how different output combination methods affect model performance. In this experiment, we use FC-KAN (DoG+BS) with several output combination methods: sum, element-wise product, addition of sum and element-wise product, quadratic and cubic function representations, concatenation, and linear transformation of the concatenated output. Inspired by the work of Altarabichi [11] on DropKAN [46], we also include the maximum, minimum, and average outputs for comparison.
@@ 4.3 Comparison of Combination Methods
Dataset Combined Method Train. Acc. Val. Acc. F1 Time (seconds)
@@
From the results in Table 3, the quadratic function best represents the combination output and outperforms other combinations, although its models require more training time. It is clear that the outputs combined by element-wise operations consistently outperform other methods, demonstrating superior accuracy. This indicates that element-wise combinations are more effective in capturing and integrating relevant features from the data, leading to better performance. Meanwhile, the output combination by concatenation shows the worst results. The linear transformation applied to the concatenated output outperforms concatenation alone, but still achieves only average performance. Similarly, the maximum, minimum, and average outputs do not deliver superior results.
In MNIST, besides the quadratic functions, the addition of sum and element-wise product demonstrates very competitive performance while requiring the least training time. Except for concatenation, maximum, and minimum, all other combinations can easily achieve 100% training accuracy. In Fashion-MNIST, the element-wise product combination is only surpassed by the quadratic function, but the plus point is it takes 6.3% less training time. The addition of sum and element-wise product achieves the best training accuracy, followed by the quadratic function and the linear transformation of concatenation.
Contrary to our expectations, the cubic function representation only achieves moderate performance. However, it takes the longest training time on MNIST and ranks just behind the quadratic function in terms of training time on Fashion-MNIST. Initially, we hypothesized that the cubic representation could capture more data features, but it appears that the excessive number of element-wise operations may hinder feature extraction, potentially leading to reduced performance. This experiment demonstrates that using higher-degree functions may not necessarily enhance model performance and can also increase computational complexity.
4.4 Model Performance with Limited Data
Rather than training FC-KAN models on the full data to determine the optimal configuration, we investigate whether using smaller portions of the data can yield comparable insights. This approach not only helps identify the best function combinations but also significantly reduces training time. Additionally, it allows for testing a greater number of configurations and provides early indications of model performance on the full training data. In this experiment, we evaluate 6 FC-KAN variants—FC-KAN (DoG+BS), FC-KAN (DoG+RBF), FC-KAN (DoG+BASE), FC-KAN (BS+RBF), FC-KAN (BS+BASE), and FC-KAN (RBF+BASE)—using 1%, 5%, and 10% of the data, with a quadratic function representation for combining function outputs. Each model is trained over 5 independent runs with the same configuration, and we calculate the average performance.
Figure 5: The validation accuracy values of the models across various data subsets.
Figure 5 illustrates the performance of the FC-KAN models across different training subsets and function combinations. In the MNIST dataset, the RBF+BASE combination demonstrates superior performance with 1%, 5%, and 10% of the data; however, the highest performance on the full training data is recorded for the DoG+BS combination. This observation underscores the challenges of predicting the optimal function combination when working with limited training data. In the Fashion-MNIST dataset, DoG+BS achieves the best performance with 1% of the data, while DoG+RBF excels with both 5% and 10%. These results suggest that DoG+BS and DoG+RBF may perform well with the full training dataset. Indeed, DoG+BS consistently outperforms other combinations in full training, with DoG+RBF serving as a strong competitor, as in Table 2. Overall, our findings indicate that accurately predicting the performance of FC variants on the complete training dataset based solely on their performance with smaller subsets is challenging. This variability may depend on the specific datasets and the portions of data used for training.
@@ 6 Conclusion
We introduced FC-KAN, which uses various popular mathematical functions to represent data features and combines their outputs using different methods, primarily through element-wise operations in low-dimensional layers, to address image classification problems. In the experiments, we designed FC-KAN to combine pairs of functions, such as B-splines, wavelets, and radial basis functions, using several output combinations on the MNIST and Fashion-MNIST datasets. We found that FC-KAN outperformed MLP and other KAN models with the same network structure, although it required more training time. Among the variants, FC-KAN (DoG+BS), which combines the Difference of Gaussians and B-splines with a quadratic function representation in the output, achieved the best results on both datasets. Our experiments also show that it is not easy to detect the function combination performance in the full training data based on small-scale experiments. FC-KAN demonstrates promising potential in using function combinations to design KANs and improve model performance. However, there are several aspects, such as exploring other functions and combinations to extract data features for further model enhancement, that we would like to investigate. These will be the focus of our future work.